OpenStudy (anonymous):

Write an equation of a line with a slope of 0 and passing through the point (5,4).

OpenStudy (anonymous):

well y=mx+b is y=0+b

OpenStudy (anonymous):

which is also 4=0+b which b is just 4 so y=4

OpenStudy (cwrw238):